OpenStudy (anonymous):

8p=12-2q q+4p=6 Which statement is true about the lines graphed from the equations ? A. they coincide B. they are parallel C.They are perpendicular . D.they intersect but are not perpendicular .

OpenStudy (anonymous):

8p=12-2q (1) q+4p=6 (2) *2 2q+8p=12 8p=12-2q equation (1)= equation (2)

OpenStudy (anonymous):

solve the given two equtions.... we will get same equtions... so we can say that they are coincide ..

OpenStudy (anonymous):

so we just solve them & that's that ? okay thanks :) @chetan552

OpenStudy (anonymous):

You're welcome...:-)